OpenStudy (spareb665):

X=-In2/(3In2+In5) round the answer 2 digits after the decimal sign.

OpenStudy (mertsj):

What is another form of 3ln2?

OpenStudy (spareb665):

In(8)?

OpenStudy (mertsj):

Yes. And what is another form of ln8 + ln5 ?

OpenStudy (spareb665):

idk In(40)?

OpenStudy (mertsj):

Yes. So now we have:

OpenStudy (mertsj):

\[x=- \frac{\ln 2}{\ln 40}\]

OpenStudy (mertsj):

So pick up your calculator and find the ln of 2, divide it by the ln of 40, round the answer to two decimal places and change it to its opposite.

OpenStudy (mertsj):

What did you get?

OpenStudy (spareb665):

I got -0.1879018247

OpenStudy (spareb665):

but im not sure if its right 2^3x+1=5^-x= (3x+1)In2=-xIn5 3xIn2+In2=-xIn5 3xIn2+xIn5=-In2 X(3In2+In5)=-In2 X=-In2/(3In2+In5) =-In2/In40 = -0.1879018247

OpenStudy (mertsj):

Round it to -.19

OpenStudy (spareb665):

did i do that right? x= -.19 ?

OpenStudy (mertsj):

Replace x with -.19 and see if it makes the original equation true.

OpenStudy (mertsj):

Also, that is what I got.